A box is pulled on level ground by a force of 40 N that makes 45° with the ground.


Find a) the force that tends to lift the box, and b) the force that tends to pull the box forward.


1
Expert's answer
2021-04-02T10:34:36-0400

"\text{expand the force F into vertical and horizontal projections}"

"\vec{F}= \vec{F}*\cos\alpha+\vec{F}*\sin\alpha"

"\text{a) the force that tends to lift the box is } \vec{F}*\cos\alpha= 40*\frac{1}{\sqrt2}\approx28.3 J"

"\text{b) the force that tends to pull the box forward is}\vec{F}*\sin\alpha= 40*\frac{1}{\sqrt2}\approx28.3 J"

Answer:a) 28.3J b) 28.3J